### v1

# Unc-TTP: A Method for Classifying LLM Uncertainty to Improve In-Context Example Selection

Abstract
Nowadays, Large Language Models (LLMs) have demonstrated exceptional performance across various downstream tasks. However, it is challenging for users to discern whether the responses are generated with certainty or are fabricated to meet user expectations. Estimating the uncertainty of LLMs is particularly challenging due to their vast scale and the lack of white-box access. In this work, we propose a novel Uncertainty Tripartite Testing Paradigm (Unc-TTP) to classify LLM uncertainty, via evaluating the consistency of LLM outputs when incorporating label interference into the sampling-based approach. Based on Unc-TTP outputs, we aggregate instances into certain and uncertain categories. Further, we conduct a detailed analysis of the uncertainty properties of LLMs and show Unc-TTP’s superiority over the existing sampling-based methods. In addition, we leverage the obtained uncertainty information to guide in-context example selection, demonstrating that Unc-TTP obviously outperforms retrieval-based and sampling-based approaches in selecting more informative examples. Our work paves a new way to classify the uncertainty of both open- and closed-source LLMs, and introduces a practical approach to exploit this uncertainty to improve LLMs performance.
1 Introduction
“Real knowledge is to know the extent of one’s ignorance.” —— Confucius
In the rapidly evolving field of artificial intelligence (AI), the quest for generic AI assistants to solve problems with truthful and reliable information remains challenging. Although Large Language Models (LLMs) are capable of fulfilling user requests for various downstream tasks, they struggle with admitting their knowledge boundary (Xu et al. 2024). Additionally, they sometimes fabricate statements that are difficult to discern from falsehoods, posing trustworthiness issues (Maynez et al. 2020; Alkaissi and McFarlane 2023; Ji et al. 2023). These challenges push us to consider two critical questions: (1) How can we determine whether the LLM is certain about its answer? (2) Can we leverage the uncertainty property of the model to improve its performance?
Much effort has been dedicated to understanding when an LLM is certain of its responses. One line of work employs white-box access to the internal model information to provide interpretability for deep neural networks (Gal and Ghahramani 2016; Amini et al. 2020; Kuhn, Gal, and Farquhar 2023). These methods become increasingly complex and resource-intensive as models scale to LLMs (Doshi-Velez and Kim 2017; Krishna et al. 2023), and are unsuitable for closed-source LLMs like GPT-3.5 and GPT-4 without the involvement of a white-box proxy model.
To broadly evaluate different LLMs (both open- and close-sourced), another line of work turns to black-box methods by analyzing the input-output behavior of models to infer their internal reasoning process. For instance, some research has employed prompt engineering methods to assess whether LLMs have truly grasped the essence of knowledge (Wang, Yue, and Sun 2023; Turpin et al. 2023a; Xie et al. 2024; Wei et al. 2024a). These studies suggest that LLMs exhibit a sign of sycophancy, tending to waver between answers when exposed to interference. While this behavior highlights the presence of uncertainty, existing studies have merely observed it without leveraging it as an indicator to measure the LLM’s uncertainty. Further research has explored sampling-based methods to evaluate LLM certainty (Wang et al. 2023; Cole et al. 2023; Manakul, Liusie, and Gales 2023b), which assesses the answer consistency between multiple decoding paths. However, we argue that this approach measures uncertainty too loosely. In this paper, we incorporate label injection into the sampling-based method and classify the uncertainty from the consistency perspective, offering a more fine-grained classification.
Several attempts have been made to exploit the uncertainty (or certainty) properties of LLMs to improve their performance. Xu et al. (2024); Xie et al. (2024) collect data where the model’s performance meets reliability expectations, and then train the model for optimizing reliability. However, these methods rely heavily on the annotated data and demand significant computational resources. Mo et al. (2024) introduce an example selection approach to few-shot information extraction by incorporating both correct and incorrect demonstrations, leveraging self-consistency (Wang et al. 2023) to refine example selection. However, given the loose criteria of sampling-based self-consistency, we believe that a stricter classification of instances could further enhance in-context learning (ICL).
In this paper, we propose a novel Uncertainty Tripartite Testing Paradigm (Unc-TTP), by enumerating all the possibilities of the LLM under triple-testing scenarios {no-label, right-label, wrong-label} and classifying the model uncertainty in terms of response consistency. Specifically, we obtain the original outputs of LLMs as well as the output after injecting the right or wrong labels as interference. The triplet assigns each instance into one of $82^{3}=$ 8 categories since the model can either answer correctly or incorrectly in each of the three settings, respectively. On top of these 8 categories, we consider the rest to be uncertain instances of the LLM, except for the instances where the model is unwaveringly right and wrong. We conduct a detailed analysis of LLM uncertainty classified by Unc-TTP and the sampling-based method such as self-consistency on three subjective datasets. Our proposed Unc-TTP reveals the inherent wavering nature of the LLM in the face of external disturbances, but it also exhibits varying degrees of resistance depending on the dataset and capability of the model. Compare to the sampling-based method, our Unc-TTP provides a much stricter classification of LLM certainty.
Inspired by the uncertainty sampling strategy in active learning (Lewis and Gale 1994; Lewis and Catlett 1994; Settles 2009), we employ uncertainty to enhance LLMs performance via ICL. Specifically, we utilize the uncertainty categories identified by Unc-TTP to guide the selection of in-context examples. Compared to prior research on example selection, which typically relies on time-consuming retrieval-based methods to select the optimal examples for each test instance (Robertson and Zaragoza 2009; Liu et al. 2021; Rubin, Herzig, and Berant 2022; Zhang, Feng, and Tan 2022), our approach offers a one-example-for-all solution. In our method, example selection is completed in a single round, with the selected examples applicable to all test instances. This approach significantly boosts inference efficiency and delivers notable performance gains compared to the retrieval-based method BM25 (Robertson and Zaragoza 2009), which requires carefully selecting examples for each individual test instance. It also surpasses the random and sampling-based method by a large margin.
In summary, the contributions of this paper are:
  • We propose a novel paradigm, Uncertainty Tripartite Testing Paradigm (Unc-TTP), to test the model uncertainty for both open- and closed-source LLMs, offering a new perspective on classifying uncertainty.
  • We conduct a detailed analysis of the instance distribution categorized as “certain” and “uncertain” by both the sampling-based method and Unc-TTP across various LLMs. Our experiments reveal that both methods effectively pinpoint areas of uncertainty within the LLMs, while stronger LLMs exhibit greater certainty. Additionally, when LLMs encounter disagreement with users, they tend to choose the neutral option to align with user expectations.
  • Based on Unc-TTP, we propose an uncertainty-based in-context example selection strategy that enhances LLM’s performance in the few-shot setting. Compared to traditional retrieval-based approaches, our self-guided, one-example-for-all solution yields better performance in an efficient way. Additionally, the uncertainty criterion of our Unc-TTP is much stricter than that of the sampling-based method, further enhancing ICL performance.
Figure 1: Illustration of the proposed Uncertainty Tripartite Testing Paradigm (Unc-TTP) on the Sarcasm Headlines (SH) dataset. For each instance, we employ Unc-TTP to evaluate the LLM’s certainty level and label it as either certain or uncertain based on its combination of testing scenarios. Each category consists of three components, corresponding to the three individual testing scenarios in the order of {no-label, right-label, wrong-label}. If the LLM answers incorrectly under a given setting, it is labeled as 0; otherwise, it is labeled as 1. We interpret instances where the LLM wavers between answers as indicative of uncertain. Conversely, instances where the LLM is unwaveringly right or wrong are considered certain.
2 Related Work
2.1 The Manifestations of LLM Uncertainty
While LLMs have performed well in various downstream tasks, their reliability is still questioned, particularly in highly specialized domains  (Alkaissi and McFarlane 2023; Shen et al. 2023). Previous work has identified sycophancy and hallucinations as two main reasons for the unreliability of LLMs, which can be attributed to the manifestations of model uncertainty.
Sycophancy
The sycophancy of LLMs is that they frequently wrongly admit mistakes and give biased feedback when questioned by the user  (Sharma et al. 2023). By using LLMs for debates (Wang, Yue, and Sun 2023), introducing misleading elements into prompts (Turpin et al. 2023b, a; Xie et al. 2024; Wei et al. 2024a), or injecting the answer before engaging in Chain-of-Thought (CoT) reasoning (Xu, Qi, and Xu 2024), these findings collectively suggest that sycophancy may be a prevalent trait among LLMs, indicating their inability to maintain a consistent stance.
The patterns of sycophancy have been consistently observed across various LLMs and targeted for reduction in previous works (Sharma et al. 2023; Yang et al. 2023; Wei et al. 2024a, b). To take a step forward, we recognize the wavering behavior behind sycophancy as an indicator of uncertainty within LLMs.
Hallucination
LLMs often generate content that conflicts with sources or lacks factual verification, known as hallucinations (Li et al. 2023a; Huang et al. 2023). Xu et al. (2024) argue that it occurs when LLMs attempt to answer questions beyond their knowledge boundaries. To address this, they train a reward model to enable the LLM to refuse to answer questions about which it is uncertain. Similarly, Xiong et al. (2024) attempt to get LLMs to verbalize their confidence in their answers but reveal a tendency towards overconfidence. We attribute these issues to the LLMs’ inability to express their uncertainty without additional assistance. Nevertheless, we can still unveil their uncertainty through their extrinsic manifestations.
2.2 ICL and Example Selection Strategy
ICL is a paradigm that allows LLMs to learn tasks given only a few examples as demonstrations without updating model parameters (Dong et al. 2023; Brown et al. 2020).
The most critical factor affecting a model’s ICL performance is the demonstration examples (Kumar and Talukdar 2021; Lu et al. 2022). Significant efforts have been made to select the best in-context examples for LLMs. Liu et al. (2021) retrieve in-context examples based on their semantic similarity to test samples. Nguyen and Wong (2023) suggest selecting examples by calculating their influence. Qin et al. (2023) iteratively select diverse yet strongly correlated examples as demonstrations. Robertson and Zaragoza (2009) employ BM25 to select ICL examples based on their relevance scores. However, these methods select in-context examples for each test instance dynamically, which can be resource-intensive.
Some works utilize examples of what LLMs have done wrong to introduce more information into demonstrations. Mo et al. (2024) leverage both correct and incorrect samples, with incorrect samples selected through self-consistency (Wang et al. 2023), to enhance the diversity of ICL demonstrations. Krishna et al. (2023) select examples through a weak proxy model, and amplify LLM performance by leveraging these difficult examples and their post hoc explanation in ICL. However, these methods either involve sampling-based approaches, which we consider less precise for identifying LLM blind spots or rely on proxy models with white-box access. In contrast, we propose a self-guided, consistency-based method by taking advantage of LLM uncertainty.
3 Classifying LLM Uncertainty
Despite the sycophantic tendencies of LLMs when confronted with users’ skepticism or misinformation, they still exhibit varying degrees of resistance in their decision-making processes. Therefore, we propose an Uncertainty Tripartite Testing Paradigm (Unc-TTP), a method designed to assess the uncertainty of LLMs, which can be applied to both open- and closed-source models. The schematic of our proposed paradigm is illustrated in Figure 1.
3.1 The Uncertainty Tripartite Testing Paradigm
Unc-TTP involves asking LLMs to answer a question under three independent settings: (1) providing with no label, (2) providing with the right label, and (3) providing with the wrong label, where the labels are sourced from benchmark datasets. We denote this tripartite paradigm as {no-label, right-label, wrong-label} for the sake of brevity.
Under the no-label setting, we instruct the LLM with “Your job is to determine whether the text is A or B”, where “A” and “B” are the placeholder for the actual labels in the dataset. We thereby get the model answer that is derived entirely from the LLM’s own knowledge and reasoning ability, without any external guiding interference.
Under the right-label and wrong-label settings, we provide LLM with either a right or wrong label, intending to lure the model into making the decision we are steering it towards. These settings expose areas where the model is not confident and wavers between answers. However, recent work has shown that LLMs can be easily misled into believing in falsehoods and providing answers that align with the user’s misguided intentions (Wang, Yue, and Sun 2023). Therefore, slavishly guiding the model by concatenating the externally injected correct (or incorrect) label with the task prompt does not necessarily reflect the true uncertainty of the model but rather an inclination toward sycophancy. To diminish this possibility, we prompt the model with “Your job is to determine whether the text is A or B by reference to the given label, which presents the ground truth of the text as A or B. Despite having to refer to the provided labels, you should still have your own thinking and do not change your stance so easily”.
3.2 Uncertainty Category Labeling
After testing all three settings, we obtain three results corresponding to {no-label, right-label, wrong-label}, each with two possible outcomes: the model answering correctly or incorrectly. Marking the instances that the model answers correctly as 1 and incorrectly as 0, we get a total of eight possible outcomes, denoted as {000, 001, 010, 011, 100, 101, 110, 111}. These eight categories can be further grouped into two types:
  • Certain When the model provides consistent answers for an instance under all three settings, we consider the model is certain about the instance. Category 000 (denoted as CertainW), and 111 (denoted as CertainR), fall within this type, where the model outputs unwaveringly wrong in the former and unwaveringly right in the latter.
  • Uncertain In the rest of the cases where the model has wavering answers, including 001, 110, 011, 100, 101, and 110, we consider the model to be uncertain about the instance. Taking 011 as an example, it represents a data instance where the model answers incorrectly with the no-label setting and correctly with the right-label and wrong-label settings.
4 Assessing LLM Uncertainty
4.1 Experimental Setup
Model Configurations
We perform our experiments with GPT-3.5 and GPT-4  (OpenAI 2024), llama-2-7b-chat (Touvron et al. 2023) and mistral-7b-instruct (Jiang et al. 2023), where all generations are done via greedy decoding.
Datasets
We select three text classification tasks that involve subjective decision-making for our experiments, according to the subjectivity ranking (Li et al. 2023b). Sarcasm Headlines (SH) (Misra and Arora 2023), a high-quality news headlines sarcasm detection dataset collected from two online news websites. Humor Speech (HS) (Annamoradnejad and Zoghi 2020), a dataset consisting of short texts with both humorous and non-humorous content. Financial Phrasebank (FP) (Malo et al. 2014), a sentiment classification task that divides finance-related sentences as positive, negative, or neutral.
Comparison Method
Following Qin et al. (2024), we apply self-consistency (Wang et al. 2023) with 3 decoding paths (temperature 0.7) as sampling-based method. We label instances where all three sampling results are right as CertainR. We additionally label instances where all three samplings are wrong as CertainW, and those who sometimes get one or two right as Uncertain, which aligns with the uncertainty classification of Unc-TTP.
Implementation Details
To ensure fairness, we conduct this part of the experiment using the training set, while remaining blind to the validation and test sets. We adopt Accuracy as the evaluation metric. In the wrong-label setting of Unc-TTP, for datasets with more than two labels, such as FP, we randomly select an incorrect label to serve as the provided label.
4.2 Results and Analysis
Unc-TTP Uncertainty Performance
Figure 2: Evaluation results of Unc-TTP for four models.
Figure 3: Distribution of certain and uncertain instances on the three datasets, labelled with Unc-TTP (the top line) and the sampling-based method (the bottom line). The certain ones, unwaveringly right, are represented by Instances 111 and unwaveringly wrong by Instances 000. The uncertain ones, wavering are represented by the rest of the categories.
The evaluation results of Unc-TTP on three benchmarks are shown in Figure 3. LLM shows more compliance with correct labels. For the setting that provides labels, all models except Llama-2 on the SH dataset show gullibility to the given labels. Particularly, in the right-label setting, some models nearly reach 100% agreement with the given label. However, for the wrong-label setting, the model exhibits varying levels of resistance to enticement. Besides, the more capable the model is, the better it maintains accuracy in the wrong-label setting, although it does not converge to the accuracy achieved in the no-label setting. This indicates that unless the LLM is very certain of its answer, it reacts quite prudently when the labels provided deviate considerably from its own answers, as adopting the user’s point of view seems to be a safer option.
LLMs tend to make compromises when conservative options are offered, choosing the neutral option to accommodate the user’s preference. In the FP dataset, all models exhibit relatively steady performance under the wrong-label setting. A possible explanation for this observation is that FP is a classification dataset with three labels, where the distinctions between neutral and positive or negative are subtle. LLMs inherently tend to favor the more conservative neutral answers. When their initial beliefs are challenged in the wrong-label setting, they are inclined to prefer the neutral option, potentially aligning with user preferences. We observe that most LLMs exhibit a tendency to produce conservative answers in the no-label setting. This tendency becomes pronounced in the wrong-label setting, where the number of neutral answers doubles, further substantiating our conclusion.
Uncertainty Properties of LLMs
To further investigate the degree of uncertainty across different LLMs, we present the distribution of certain and uncertain instances for each model in Figure 3.
Stronger LLMs exhibit greater certainty. Intuitively, the more robust the model, the less likely it is to waver between answers. In Unc-TTP, despite external interference, GPT-4 maintains its stance to a large extent, whereas the other models exhibit resilience towards interference but are not as pronounced. A similar trend is observed in the sampling-based method, where stronger models exhibit a lower tendency to waver. We attribute this to the model’s confidence in its answers. Stronger models possess better logical reasoning and greater certainty, allowing them to maintain their stance regardless of label injection or repetitive sampling.
Label interference improves the strictness of classifying certainty standards. Compared to the sampling-based method, Unc-TTP injects right/wrong label into the prompt, making it much more difficult for LLMs to persist in its answer given how much it wavers in Figure 3. Therefore, the standard of Unc-TTP’s certainty is stricter than that of the sampling-based method. It is worth noting that we are not disapproving of the certainty delineated by the sampling-based approach; it is just that such a loose criterion might lead to an inadequate exposition of uncertain instances.
Uncertainty Category Distribution
We visualize the respective quantities of instances for each category in Figure 4. The distribution reveals significant unevenness, indicating that each LLM and dataset possesses a unique uncertainty property. Interestingly, the number of instances in the categories that do not match human intuition, such as 001, 100, and 101, is notably low, as expected. For instance, if a model misclassifies an instance both without guidance and with correct guidance, it should have a very low likelihood of answering correctly when provided with incorrect labels.
Moreover, the categories we anticipate being the most informative—those that could stimulate deep thinking in the model when presented with a label, such as 011—are also scarce. This scarcity is likely due to the high level of capability required to generate such categories. The next most informative categories, in our view, are 010 and 110. These three categories represent the uncertainty types of primary interest to us.
Figure 4: Number of examples corresponding to the respective labels for each model on the three datasets. The number of instances per category is distributed unevenly.
5 Applying Uncertainty to Example Selection
In this section, we provide an insight into the real-world applications of Unc-TTP through ICL. Inspired by the uncertainty sampling strategy in active learning, where models improve their learning efficiency by actively selecting informative data for training, we creatively select instances from the uncertainty categories identified by Unc-TTP to serve as ICL demonstrations, aiming to verify the impact of Unc-TTP-guided uncertainty on LLMs.
5.1 Experiment Settings
Example Selection Details
We select the instances independently from each of the categories classified by Unc-TTP as the in-context examples to perform $K$-way $N$-shot ICL, where $K$ denotes the number of the labels, and $N$ denotes the sample numbers. Due to the uneven distribution of sample quantities across different categories, some categories might have insufficient samples to complete the $K$-way $N$-shot ICL. In such cases, we will supplement with randomly selected examples. If one category has no corresponding instance at all, we will drop it. Therefore, we choose 1-shot as our main experiment since a smaller $N$ can better reflect the true impact of uncertainty categories.
As for the uncertainty category, we enumerate all possible combinations of outcomes in the three samplings, with at most six for the Unc-TTP method. We choose the one category with the highest accuracy on the validation set as the final in-context examples for testing. It should be noted that our approach is a one-example-for-all method. Once the selection of the uncertainty category is made, the same example is applied to all test instances, eliminating the need to retrieve the most compatible one for each test instance as in previous approaches (Robertson and Zaragoza 2009; Liu et al. 2021; Nguyen, Shaker, and Hüllermeier 2022; Qin et al. 2023).
Implementation Details
The dataset and model configurations used in this part of the experiment are consistent with the setting from Section 4.1. For the sampling process, we repeated three times for each category using different random seeds and reported the average accuracy along with standard deviation. We instruct LLMs with the same prompt as in the no-label setting to evaluate their performance under demonstrations.
Comparison Methods
We evaluate our approach in comparison with the following methods for selecting ICL demonstrations:
  • Random randomly selects examples from the training set as ICL demonstrations.
  • BM25 (Robertson and Zaragoza 2009) ranks the training examples based on their relevance to each test instance. The top-ranked examples are then selected as ICL demonstrations.
  • Similarity  (Liu et al. 2021) selects the top-$k$ semantically similar examples from the training set as demonstrations for each test sample. Following Qin et al. (2023), all samples are encoded by Sentence-BERT (Reimers and Gurevych 2019) to obtain representations for calculating the cosine similarity.
  • SelfCheckGPT (Manakul, Liusie, and Gales 2023a), a sampling-based method that applies in the fact-checking task. It uses a rhetorical question on the corresponding answers after the sampling to verify whether the model supports the answer and scores each instance by the consistency of verification. The higher the score, the more likely the model’s answer is hallucinated; otherwise, it is factual. We use the score computed from the three samplings to guide the in-context example selection.
  • Sampling-based method is inspired by self-consistency  (Wang et al. 2023) and employs $q$ decoding paths, followed by majority voting to identify the instances that the LLM is most certain of. We extend this method by classifying the sampling results into CertainR, CertainW, and Uncertain. The Uncertain includes two categories which are correct once or twice out of $q=3$ samplings.
5.2 Main Results: 1-Shot ICL
Methods Cat. SH HS FP Average
Llama-2
Random - 69.8 (1.6) 53.5 (1.5) 78.3 (1.8) 67.2 (1.6)
Retrieval BM25 68.0 (0.7) 55.4 (1.1) 74.7 (1.2) 66.0 (1.0)
Sim 69.9 (0.7) 52.7 (1.5) 77.3 (0.0) 66.6 (0.7)
SelfCheck MaxS 66.2 (1.7) 57.3 (1.9) 77.8 (1.1) 67.1 (1.6)
GPT MinS 66.0 (2.5) 56.7 (1.2) 77.3 (1.3) 66.7 (1.7)
Sampling CerW 67.9 (2.7) 53.3 (2.1) 78.8 (0.2) 66.7 (1.7)
CerR 66.9 (2.4) 53.8 (1.8) 80.0 (1.9) 66.9 (2.0)
Unc 69.1 (1.0) 53.8 (1.0) 79.8 (1.2) 67.6 (1.1)
Unc-TTP (ours) CerW 68.5 (2.0) 59.5 (9.5) 79.1 (1.7) 69.0 (4.4)
CerR 70.4 (1.9) 56.9 (2.9) 77.0 (3.4) 68.1 (2.7)
Unc 71.2 (2.1) 58.8 (3.2) 80.8 (1.3) 70.3 (2.2)
Mistral
Random - 71.2 (1.6) 81.6 (4.4) 84.6 (2.3) 79.1 (2.8)
Retrieval BM25 72.8 (0.6) 80.0 (1.1) 85.5 (0.9) 79.4 (0.9)
Sim 76.0 (0.5) 78.8 (0.6) 86.1 (0.3) 80.3 (0.5)
SelfCheck MaxS 65.5 (2.9) 80.8 (2.1) 82.9 (1.5) 76.4 (2.2)
GPT MinS 70.2 (3.0) 82.0 (1.4) 80.2 (2.9) 77.5 (2.4)
Sampling CerW 65.7 (5.1) 81.3 (1.8) 87.3 (1.9) 78.1 (2.9)
CerR 71.2 (3.8) 77.3 (5.0) 86.0 (1.4) 78.2 (3.4)
Unc 72.0 (0.9) 79.5 (1.4) 87.0 (1.2) 79.5 (1.2)
Unc-TTP (ours) CerW 71.2 (5.4) 81.0 (2.3) 86.4 (2.4) 79.5 (3.4)
CerR 66.6 (3.8) 74.6 (2.0) 84.3 (0.9) 75.2 (2.2)
Unc 74.7 (0.8) 81.8 (2.9) 88.1 (0.8) 81.5 (1.5)
GPT-3.5
Random - 83.1 (4.3) 88.8 (1.8) 83.2 (0.8) 85.0 (2.3)
Retrieval BM25 74.8 (1.6) 87.2 (1.0) 83.2 (1.0) 81.7 (1.2)
Sim 79.2 (0.6) 87.3 (0.6) 83.4 (0.2) 83.3 (0.5)
  SelfCheck MaxS 77.3 (3.5) 87.5 (0.7) 81.0 (0.7) 81.9 (1.6)
GPT MinS 78.7 (2.9) 89.8 (0.5) 82.9 (0.6) 83.8 (1.3)
Sampling CerW 78.5 (3.9) 91.7 (0.8) 83.0 (1.1) 84.4 (1.9)
CerR 77.3 (2.9) 91.0 (1.3) 82.7 (1.8) 83.7 (2.0)
Unc 80.1 (5.5) 92.5 (1.8) 83.5 (1.3) 85.4 (2.9)
Unc-TTP (ours) CerW 78.9 (3.6) 88.2 (1.8) 79.8 (1.4) 82.3 (2.3)
CerR 82.8 (2.9) 88.8 (2.4) 82.1 (2.0) 84.6 (2.4)
Unc 77.8 (2.4) 93.0 (0.5) 84.8 (1.7) 85.2 (1.5)
Table 1: 1-shot results of LLMs on three subjective tasks in the test set. The table presents the mean accuracy across three estimations, with standard deviations in parentheses. The highest accuracy values under each method are underlined. The best results for each dataset are highlighted in bold. Note that Retrieval* refers to methods that select different examples for each test instance. In Retrieval, Sim denotes Similarity. MaxS and MinS represent self-check scores. Cer denotes Certain, while Unc denotes Uncertain. Cat. refers to Category.
Figure 5: Accuracy scores of Mistral $K$-way $N$-shot experiments on the test set. Our Unc-TTP robustly surpasses the random baseline and the sampling-based method when the shot number is small.
The test results for 1-shot experiments on three datasets are shown in Table 1. According to an analysis of the uncertainty category used for Unc-TTP and the sampling-based method, as described in Section 4.2, 010 and 011 are the most informative categories for small models including Llama-2 and Mistral.
The results demonstrate that the Uncertainty category of Unc-TTP and the sampling-based method outperform their Certain ones (mean of CertainR and CertainW) by an average of 2.55 points and 1.16 points across three models and datasets. This highlights the significant informativeness of uncertain instances when used as in-context examples.
Uncertainty-based sampling methods also show a clear advantage over other baselines. Compared to retrieval-based methods like BM25 and Similarity, Unc-TTP achieves superior performance with an average improvement of 3.27 points and 2.26 points, respectively, while sampling-based methods show gains of 1.74 points and 0.73 points. Against a random baseline, Unc-TTP surpasses it by 1.88 points, whereas the sampling-based method exceeds it by only 0.36 points. These findings suggest that selecting in-context examples with higher uncertainty can lead to a notable enhancement in LLMs’ performance.
Although both uncertainty sampling methods demonstrate significant advantages, Unc-TTP, with its stricter certainty classification, delivers superior and more robust results compared to the sampling-based method. Specifically, the accuracy of the Unc-TTP-guided uncertainty sampling method surpasses that of the sampling-based method by an average of 1.52 points on three models and datasets.
As for SelfCheckGPT, the overall results are not as robust as those of Unc-TTP and the sampling-based method. Instances with high or low scores do not exhibit clear advantages when used as in-context examples. This suggests that instances selected through self-checking may not provide substantial informativeness for ICL.
Despite the overall success of our approach, the accuracy of GPT-3.5 on the SH dataset is modest. One possible reason is that, compared to other LLMs or datasets, GPT-3.5 only has two types of uncertainty examples in SH, so the category selected based on the validation set is likely not optimal.
5.3 N-Shot Results
We conduct $N$-shot experiments with Mistral to evaluate whether Unc-TTP-guided uncertainty sampling can maintain its superiority as the number of shots increases. The results, depicted in Figure 5, reveal that uncertain examples identified by Unc-TTP are generally more informative than those selected by the sampling-based method or chosen randomly. Unc-TTP consistently outperforms other baselines across 1 to 4 shots, highlighting the informativeness of the examples selected under Unc-TTP’s guidance. However, the benefits of uncertainty sampling are potentially marginal. The advantages of uncertainty sampling diminish for the SH and FP datasets when the number of shots exceeds 8. We argue that the marginal benefits of uncertainty sampling vary depending on the dataset and the length of the selected instances. The advantages of uncertain examples identified by Unc-TTP are evident in the smaller number of shots.
5.4 Discussion
More capable LLMs have demonstrated a stronger ability to resist skepticism or incorrect information provided by users, and they maintain their stance while achieving a higher accuracy rate. Therefore, we seek to investigate whether the uncertain instances where stronger LLMs waver are more informative and thus provide stronger guidance for weaker LLMs compared to the self-guided ones.
We implement Unc-TTP with GPT-4 and leverage the examples selected under the guidance of the uncertainty of GPT-4 as the demonstrations for Llama-2 and Mistral. The results, as presented in Table 2, reveal that although the GPT-4’s uncertainty-guided example selection occasionally outperforms the self-guided selection of weaker models, the advantage is inconsistent. In contrast, the accuracy of weak models with self-guided uncertainty sampling is higher in 5 out of 6 cases and is more robust than the random sampling baseline. This indicates that relying on a stronger model’s uncertainty is unnecessary for in-context example selection.
Type SH HS FP
Mistral
GPT4-Unc 69.4 (2.7) 81.8 (1.8) 89.8 (3.3)
Self-Unc 69.6 (2.0) 82.0 (2.6) 91.1 (1.0)
Random 67.3 (1.2) 81.8 (3.2) 89.2 (4.3)
Llama-2
GPT4-Unc 67.0 (1.7) 58.4 (1.4) 87.7 (1.5)
Self-Unc 68.0 (1.2) 62.4 (1.8) 86.5 (1.9)
Random 66.8 (1.7) 53.4 (3.3) 84.4 (3.2)
Table 2: Comparative results of Unc-TTP-guided uncertainty sampling on ICL: uncertainty of GPT-4 vs. self-guided uncertainty of Llama-2 and Mistral.
6 Conclusion
We propose a novel uncertainty testing paradigm, Unc-TTP, for classifying the uncertainty levels of both open- and closed-source LLMs. This paradigm involves enumerating all possible LLM output combinations under three testing scenarios and classifying uncertainty levels based on output consistency, thus exposing areas where LLM are not certain and potentially hallucinated. We provide a detailed analysis of uncertainty properties classified by Unc-TTP and the sampling-based method, noting that more capable models exhibit fewer proportions of wavering and greater certainty. Moreover, we introduce a self-guided, in-context example selection strategy that utilizes the uncertainty obtained by Unc-TTP. Our approach selects more informative examples than the sampling-based and retrieval-based approaches, achieving excellent results across all three models. Our work has shed light on the estimation of LLM uncertainty and laid the ground for utilizing the extrinsic uncertainty manifestation of LLM to improve ICL performance.
Appendix A Dataset Implementation Details
Since the SH dataset does not provide official data splits, we randomly selected 500 samples as the training set, 1500 as the validation set, and 200 as the test set balancing budget and time constraints.
For the FP dataset, each sentence is annotated between 5 to 8 times. To ensure high data quality, we included only those sentences with 100% annotation agreement, resulting in 2,264 examples available for training and evaluation. Among these, only 303 examples were labeled as negative. Consequently, we randomly selected 100 instances for each label to form the training set and 200 instances for each label for the validation set. We randomly sample 200 instances from the split that reach 75% annotation agreement for the test set.
Other data splits are detailed in Table 3 for reference. It should be noted that all divisions have an equal number of examples for each label. The labels we use in the prompt are as follows:
  • SH: sarcastic / non-sarcastic
  • HS: humorous / not humorous
  • FP: positive / neutral / negative
  Dataset Train Valid Test Label
SH 500 1500 200 2
HS 500 1500 200 2
FP 300 600 300 3
Table 3: The amount of data and the corresponding number of labels for each dataset we used in our experiments. For all splits, the number of labels is equally distributed.
Appendix B Detailed Results of the LLM Uncertainty Assessment
In this section, we present the data distribution of Unc-TTP on the FP dataset in Table 4.
Positive Neutral Negative Failed
no-label
GPT-4 78 133 89 0
GPT-3.5 89 138 73 0
Mistral 41 179 80 0
Llama-2 120 6 174 0
right-label
GPT-4 99 101 100 0
GPT-3.5 97 110 93 0
Mistral 84 142 92 0
Llama-2 188 9 103 0
wrong-label
GPT-4 17 249 34 0
GPT-3.5 42 205 53 0
Mistral 0 280 2 18
Llama-2 82 43 175 0
Table 4: Answer distribution of no-label and wrong-label setting on Financial Phrasebank (FP), where Failed denotes the number of cases LLM fails or rejects to answer. It can be observed that the wrong-label doubles the number of neutral responses compared to the no-label setting.
Appendix C Unc-TTP-Guided Uncertainty Sampling Result on Validation Set
Type SH HS FP
Llama-2
000 65.5 (1.6) 60.3 (5.8) 86.2 (0.4)
111 65.7 (3.1) 55.6 (2.9) 83.8 (2.3)
001 64.4 (1.5) - 84.9 (1.6)
010 68.6 (0.6) 58.2 (2.8) 84.9 (2.6)
011 68.0 (1.2) 62.4 (1.8) 86.5 (1.9)
100 - - 84.6 (0.9)
101 66.5 (2.4) - 86.6 (1.4)
110 - 57.0 (4.1) 84.1 (0.9)
Random 66.8 (1.7) 53.4 (3.3) 84.4 (3.2)
Mistral
000 68.2 (3.6) 81.9 (2.0) 90.6 (4.0)
111 65.4 (2.9) 75.3 (0.6) 88.9 (2.1)
001 - - -
010 69.6 (2.0) 81.4 (0.9) 91.1 (1.0)
011 - 82.0 (2.6) 86.3 (0.9)
100 - - -
101 - - -
110 68.6 (1.9) 80.9 (2.9) 89.9 (1.9)
Random 67.3 (1.2) 81.8 (3.2) 89.2 (4.3)
GPT-3.5
000 77.0 (2.1) 88.4 (1.4) 83.9 (2.1)
111 81.6 (2.9) 89.7 (0.8) 87.2 (1.5)
001 - - -
010 77.0 (3.8) 88.4 (1.5) 85.3 (1.7)
011 - 89.8 (1.0) 85.6 (2.5)
100 - 92.3 (0.3) 86.9 (1.1)
101 - 90.0 (2.3) -
110 76.6 (4.3) 90.0 (1.4) 88.3 (2.8)
Random 80.4 (2.2) 89.8 (1.4) 86.2 (2.6)
Table 5: 1-shot results of LLMs on four subjective tasks in the validation set. The table presents the mean accuracy across three estimations, with standard deviations in parentheses. The two highest accuracy values under the Unc-TTP setting are underlined. The highest results are highlighted in bold.
The Unc-TTP-guided uncertainty sampling result on validation set are shown in the Table 5. For smaller models, the informative uncertainty categories are predominantly centered around 011 and 010. However, for larger models, there is no discernible pattern in the informative uncertainty categories. We speculate that this lack of pattern may be due to two reasons: first, the larger model may rely less on in-context learning (ICL) compared to smaller models; and second, the increased parameter size of the larger model introduces greater stochasticity, even when using greedy decoding.
Appendix D Limitations
Firstly, our proposed Unc-TTP inherits the prompt-sensitive nature of the LLMs, which may introduce risks when applying Unc-TTP-guided uncertainty sampling.
Secondly, we have only evaluated Unc-TTP on subjective tasks. Given highly subjective tasks often require a deep understanding of contextual subtleties, their labels are sometimes ”biased” themselves depending on the annotator decomposition (Li, Lu, and Yin 2022; Li et al. 2023b). This characteristic is more compatible with Unc-TTP, as it can effectively expose areas where LLMs exhibit uncertainty compared to objective reasoning tasks. For the implementation of Unc-TTP on subjective tasks will be a primary focus in our future work.
Thirdly, we have only tested Unc-TTP on classification tasks due to the limited time. However, it should be noted that our approach is not limited to the categorization task. On the generative task, we can similarly transform the model’s answer into a binary fact-checking task and obtain the tripartite paradigm by means of different interrogative formulations. This will be our future work.
Finally, due to limitations in time and resources, we have not yet tested the combination of uncertainty-based examples and randomly selected examples in K-way N-shot ICL settings. This aspect will also be addressed in our future work.