### v4

# Unlocking the Power of LLM Uncertainty for Active In-Context Example Selection

Abstract
Large Language Models (LLMs) have shown remarkable performance across a wide range of downstream tasks. However, it is challenging for users to discern whether the responses of LLM are generated with certainty or are fabricated to meet user expectations. In this paper, we introduce Uncertainty Tripartite Testing Paradigm (Unc-TTP), a novel method for classifying LLM uncertainty by leveraging output inconsistency. Specifically, Unc-TTP performs three rounds of sampling under varying label injection interference, enumerating all possible outcomes, and uses the degree of output inconsistency as the indicator of the LLM’s intrinsic uncertainty. To validate the effectiveness of this inconsistency-defined uncertainty, we draw inspiration from Active Learning, comparing the informativeness of actively selected in-context examples. Our experiments show that uncertainty examples selected via Unc-TTP are more informative than certainty examples. Furthermore, the Unc-TTP-guided uncertainty-based active example selection strategy outperforms existing methods, highlighting its effectiveness in classifying LLM uncertainty and enhancing in-context learning. This work not only underscores the potential of inconsistency-based uncertainty classification for both open- and closed-source LLMs but also presents a practical approach for leveraging uncertainty to improve LLM performance in real-world tasks.
Unlocking the Power of LLM Uncertainty for
Active In-Context Example Selection
Hsiu-Yuan Huang1,2, Zichen Wu1,2, Yutong Yang1,2, Junzhao Zhang1,3, Yunfang Wu1,2thanks:   Corresponding author. 1National Key Laboratory for Multimedia Information Processing, Peking University 2School of Computer Science, Peking University, Beijing, China 3School of Software and Microelectronics, Peking University, Beijing, China {huang.hsiuyuan, zhangjunzhao}@stu.pku.edu.cn, {wuzichen, yytpku, wuyf}@pku.edu.cn,
1 Introduction
"Real knowledge is to know the extent of one’s ignorance." —— Confucius
In the rapidly evolving field of artificial intelligence (AI), developing generic AI assistants that provide truthful and reliable information remains a challenge. Although Large Language Models (LLMs) are capable of fulfilling user requests for various downstream tasks, they struggle with admitting their knowledge boundary (Xu et al., 2024a). Additionally, they sometimes fabricate statements that are difficult to discern from falsehoods, causing trustworthiness issues (Maynez et al., 2020; Alkaissi and McFarlane, 2023; Ji et al., 2023), especially when confronted with users’ skepticism or misinformation (Sharma et al., 2023). These challenges push us to consider three critical questions: (1) How can we assess the uncertainty of an LLM’s response through its external behavior? (2) How can we validate the effectiveness of the assessed uncertainty? (3) Can we utilize the assessed uncertainty to enhance the performance of LLMs?
Significant efforts have been devoted to investigating the external input-output behavior of LLM to assess their underlying reasoning abilities and conjecture whether LLMs have truly grasped the essence of knowledge (Wang et al., 2023a; Turpin et al., 2023a; Xie et al., 2024; Wei et al., 2024a). These studies indicate that LLMs display some tendencies of sycophancy, often wavering between answers when confronted with human interference. While current works are still trying to find ways to eliminate this sycophantic behavior, our approach takes a different direction. We capitalize on this sycophancy to develop a method for classifying uncertainty, leveraging the identified uncertainty to enhance LLM performance.
We begin with a key observation from a simple experiment in which we sample LLM responses three times for the exact same instance, under slightly different conditions—specifically, injecting no (w/o), right, or wrong label into the prompt (Section 3.1). Surprisingly, we found that although LLMs exhibit a clear tendency of sycophancy when faced with label injection, they still show varying degrees of resistance to this behavior. In some cases, LLMs consistently maintain their stance across all three samplings, with stronger models holding their answers more firmly. Building on this observation, we hypothesize that the LLM output inconsistency can serve as an indicator of its intrinsic uncertainty.
Accordingly, we propose a fine-grained Uncertainty Tripartite Testing Paradigm (Unc-TTP), which enumerates all possible outcomes of the LLM under triple-testing scenarios {no-label, right-label, wrong-label} and classifies LLM uncertainty levels based on the response consistency. Consequently, the Unc-TTP triplet assigns each instance to one of $82^{3}=$ 8 categories, as the LLM can either provide a correct or incorrect answer in each of the three settings. Given these eight categories, we classify any instance, where the LLM answers correct or incorrect waveringly, to be uncertain.
Basically, our proposed Unc-TTP is a kind of inconsistency-defined uncertainty. Drawing inspiration from the uncertainty sampling strategy in Active Learning (AL)  (Lewis and Gale, 1994; Lewis and Catlett, 1994; Settles, 2009), we apply Unc-TTP’s classified uncertainty for active in-context example selection in ICL experiments. The preliminary study demonstrates that inconsistency-defined uncertainty examples are more informative than certainty-based ones, providing strong evidence for its potential for boosting LLM performance.
Furthermore, based on three subjective text classification tasks, we conduct experiments by applying other existing strategies for active in-context example selection. Our Unc-TTP-selected uncertainty examples significantly outperform the previous best strategy, Similarity Margatina et al. (2023), with average accuracy improvements of 3.7%, 1.2%, and 1.9% for Llama-2, Mistral, and GPT-3.5, respectively, showcasing the effectiveness of our approach in enhancing ICL performance. We also explore the scalability and transferability of Unc-TTP classified uncertainty in selecting in-context examples, demonstrating its robustness and providing compelling evidence that uncertainty defined by Unc-TTP, which is calculated based on the LLM output inconsistency, really reveals the intrinsic uncertainty of LLMs.
In summary, the contributions of this paper are:
  • To the best of our knowledge, we are the first to experimentally demonstrate that output inconsistency can act as an indicator of the LLMs’ intrinsic uncertainty via the perspective of active learning.
  • We propose a novel paradigm, Uncertainty Tripartite Testing Paradigm (Unc-TTP), to test the LLM uncertainty for both open- and closed-source LLMs, offering a new insight on classifying uncertainty.
  • Taking advantage of Unc-TTP, we propose an uncertainty-based active in-context example selection strategy, which yields greater performance gains than previous approaches in ICL.
2 Related Work
2.1 Uncertainty Behavior of LLM
While LLMs have performed well across a variety of downstream tasks, their reliability remains a topic of concern, especially in highly specialized domains such as medicine (Alkaissi and McFarlane, 2023; Shen et al., 2023). This is particularly evident when LLMs are questioned by the user and they frequently wrongly admit mistakes and give biased feedback (Sharma et al., 2023). The phenomenon where a model seeks human approval in undesirable or inappropriate ways is referred to as sycophancy (Perez et al., 2022).
Recent studies have highlighted the prevalence of sycophantic behavior in LLMs, with evidence found in various scenarios such as debates (Wang et al., 2023a), prompt manipulation to introduce misleading elements (Turpin et al., 2023a, b; Xie et al., 2024; Wei et al., 2024a), and preemptively injecting answers before engaging in Chain-of-Thought (CoT) reasoning (Xu et al., 2024b). These findings collectively suggest that sycophancy may be a prevalent trait among LLMs, indicating their inability to disregard user preference and maintain a consistent stance.
While others attempt to mitigate the sycophancy of LLM (Sharma et al., 2023; Yang et al., 2023; Wei et al., 2024a, b), our research proposes that the wavering behavior behind sycophancy can reflect the intrinsic uncertainty of LLMs. This uncertainty, in turn, could serve as a useful indicator for performing active in-context example selection.
2.2 ICL and Example Selection Strategy
ICL is a paradigm that allows LLMs to learn tasks given only a few examples as demonstrations without updating model parameters (Brown et al., 2020; Dong et al., 2023). In ICL, the quality of the demonstration examples is the most critical factor affecting LLM performance (Kumar and Talukdar, 2021; Lu et al., 2022).
Research on selecting optimal in-context examples for LLMs has been an ongoing area of study. Margatina et al. (2023) draws inspiration from the most common AL strategies, including uncertainty sampling (Gonen et al., 2024), diversity sampling (Yu et al., 2023), and similarity-based selection (Liu et al., 2021), to identify the most informative examples. However, these approaches fail to account for the unique characteristics of LLMs, particularly the intrinsic uncertainty revealed through sycophantic behavior. As we demonstrate in Section 5.1, this wavering behavior provides a more effective indicator for selecting in-context examples compared to traditional AL methods.
Other works have explored the use of incorrect examples, aiming to introduce more diverse information into the demonstrations. For instance, Mo et al. (2024) leverages incorrect examples selected through self-consistency (Wang et al., 2023b) to enhance diversity in examples. Krishna et al. (2023) selects incorrect examples using a weak proxy model and generates post-hoc explanations for these difficult cases. However, these methods either overlook the value of inconsistency—an important aspect that can help identify LLM blind spots—or depend on proxy models with white-box access.
In this paper, we propose a self-guided, uncertainty-based in-context example selection method that outperforms existing active ICL baselines.
3 Preliminary and Methodology
Despite the sycophantic tendencies of LLMs when confronted with users’ skepticism or misinformation, can they disregard users’ potentially bias information and maintain their stance? To investigate this question, we start with a simple test, whose observations lead to the development of a new methodology of classifying LLM uncertainty to improve in-context example selection: the Uncertainty Tripartite Testing Paradigm (Unc-TTP).
3.1 A Motivating Observation
The simple test is conducted under three settings: We prompt LLMs with regular query, query with right label and query with wrong label, using the prompt shown in Table 1, the LLMs and datasets described in Sections 4.1 and 4.2. The accuracy of four models is shown in Figure 2. Surprisingly, we observed varying degrees of resistance to sycophantic behavior in LLMs during their decision-making processes. In some cases, even when provided with right or wrong labels interference, the LLMs maintained their stance unwaveringly, revealing that, in some particular instances, LLMs can firmly maintain their stance across all three settings. This behavior might reflect their intrinsic uncertainty and certainty.
To further investigate this potential uncertainty (define by output inconsistency) across different LLMs, we present the distribution of wavering and unwavering instances for each model, selected using the settings described above (top), alongside the results from the vanilla method, which samples outputs three times under the no label injection setting using temperature sampling (bottom), as shown in Figure 2.
Derived from the results of Figure 2, we make the following observations. Stronger LLMs exhibit greater output consistency, with a larger percentage of them being able to unwaveringly maintain their answers across three samplings. We attribute this to the LLM’s confidence in its answers. Stronger models possess better logical reasoning ability, allowing them to maintain their stance regardless of label injection or temperature sampling. Moreover, compared to the vanilla sampling-based method, injecting right/wrong label into the prompt, making it much more difficult for LLMs to persist in its answer, therefore improves the strictness of identifying the actual instances where LLMs are certain.
Building on these key observations, we hypothesize that the LLM output inconsistency can serve as an indicator of its uncertainty. In order to prove the validity of this hypothesis, we draw inspiration from uncertainty sampling in AL and adopt AL principle for in-context example selection, where we select the most informative instances with our inconsistency-defined uncertainty as selection strategy. Since LLMs exhibit varying behaviors for different injected labels (see Appendix A), this underscores the need for a fine-grained uncertainty classification method. To address this, we propose Unc-TTP.
Figure 1: Accuracy of four models with and without label injection. See Appendix A for a detailed analysis.
(a) gpt-4
w/ label injection
(b) gpt-3.5
w/ label injection
(c) llama-2
w/ label injection
(d) mistral
w/ label injection
(e) gpt-4-vanilla
(f) gpt-3.5-vanilla
(g) llama-2-vanilla
(h) mistral-vanilla
Figure 2: Distribution of instances where the LLM remains unwavering (certain) and wavering (uncertain) across three rounds of sampling, with {no (w/o), right, wrong} label injection (top) and without label injection, using only vanilla temperature sampling (bottom), across three datasets.
Prompt
no label setting (w/o label injection) Your job is to determine whether the text is {A} or {B}.
right / wrong label setting
Your job is to determine whether the text is {A} or {B} by reference to the given label, which presents the ground truth of the text as {A} or {B}. Despite having to refer to the provided labels, you should still have your own thinking and do not change your stance so easily.
Table 1: Prompts used in Unc-TTP, where {A} and {B} denote the actual label from the training data.
3.2 The Uncertainty Tripartite Testing Paradigm
Figure 3: Illustration of the proposed Uncertainty Tripartite Testing Paradigm (Unc-TTP) on the Sarcasm Headlines (SH) dataset. For each instance, we employ Unc-TTP to evaluate the LLM’s certainty level and label it as either certain or uncertain based on its combination of testing scenarios. Each category consists of three components, corresponding to the three individual testing scenarios in the order of {no-label, right-label, wrong-label}. If the LLM answers incorrectly under a given setting, it is labeled as 0; otherwise, it is labeled as 1. We interpret instances where the LLM wavers between answers as indicative of uncertain. Conversely, instances where the LLM is unwaveringly right or wrong are considered certain.
The schematic of our proposed Unc-TTP is illustrated in Figure 3. It involves asking LLMs to answer a question under three independent settings: (1) providing with no (w/o) label, (2) providing with the right label, and (3) providing with the wrong label, where the labels are sourced from benchmark datasets. We denote this tripartite paradigm as {no-label, right-label, wrong-label} for the sake of brevity.
At the same time, to encourage the model to think independently, we prompt the model to keep its own stances when facing label interference, and the specific prompt is shown in Table 1.
The rationale behind no-label setting is that we want to get the model answer that is derived entirely from the LLM’s own knowledge and reasoning ability, without any external guiding interference. As for the right-label and wrong-label settings, we provide LLM with either a right or wrong label, intending to lure the model into making the decision we are steering it towards. These settings expose areas where the model is not certain.
Uncertainty Category Labeling
After testing all three settings, for each instance, we obtain three results corresponding to {no-label, right-label, wrong-label}, each with two possible outcomes: the model answering correctly or incorrectly. Marking the instances that the model answers correctly as 1 and incorrectly as 0, we get a total of eight possible outcomes, denoted as {000, 001, 010, 011, 100, 101, 110, 111}. These eight categories can be further grouped into two types:
Certain: When the model provides consistent answers for an instance under all three settings, we consider the model to be certain with respect to the instance. The categories 000 (denoted as CerW), and 111 (denoted as CerR) fall into this type, where the model’s outputs are unwaveringly wrong in the former and unwaveringly right in the latter.
Uncertain: In the cases where the model provides wavering answers, including 001, 110, 011, 100, 101, and 110, we consider the model to be uncertain about the instance. Taking 011 as an example, it represents a data instance where the model answers incorrectly with the no-label setting and correctly with the right-label and wrong-label settings. We denote these instances as Uncertain (Unc).
3.3 Example Selection with Unc-TTP
Taking advantage of Unc-TTP, we assign each instance in the training set with a specific uncertainty categories, and the category distribution is shown in Figure 6. The distribution reveals significant unevenness, indicating that each LLM and dataset possesses a unique uncertainty property. In this section, we select examples based on the fine-grained uncertainty categories classified by Unc-TTP.
We select the instances independently from each of the categories classified by Unc-TTP as the in-context examples to perform $K$-way $N$-shot ICL, where $K$ denotes the number of the labels, and $N$ denotes the sample numbers. Due to the uneven distribution of sample quantities across different categories, some categories might have insufficient samples to complete the $K$-way $N$-shot ICL. In such cases, we will supplement with randomly selected examples. If one category has no corresponding instance at all, we will drop it. Therefore, we choose 1-shot as our main experiment since a smaller $N$ can better reflect the true impact of uncertainty categories.
4 Experimamtal Setup
Figure 4: The relative improvement in accuracy of inconsistency-defined uncertainty examples compared to certainty examples in ICL for each model. The uncertainty examples from sampling-based methods (the bottom 2 rows) demonstrate greater robustness than those from verification-based methods (the top 2 rows), where positive accuracy gains (highlighted in red) are observed more consistently than negative ones (highlighted in blue).
4.1 Model Configurations
We perform our experiments with GPT-3.5 and GPT-4  [^1] Our experiments were done with gpt-3.5-turbo-1106 and the gpt-4-0613 checkpoint. The results for certain examples may change due to model updates.  (OpenAI, 2024), llama-2-7b-chat (Touvron et al., 2023) and mistral-7b-instruct (Jiang et al., 2023), a model that significantly outperforms the llama-2-13b-chat  (Jiang et al., 2023). All generations are done under the temperature of 0.7 except Unc-TTP example selection is done via greedy decoding.
4.2 Datasets
We select three representative subjective text classification tasks for our experiments, according to the subjectivity ranking Li et al. (2023). Sarcasm Headlines (SH) (Misra and Arora, 2023), a high-quality news headlines sarcasm detection dataset. Humor Speech (HS) (Annamoradnejad and Zoghi, 2020), a dataset consisting of short texts with both humorous and non-humorous content. Financial Phrasebank (FP) (Malo et al., 2014), a sentiment classification task that divides finance-related sentences as positive, negative, or neutral. Please refer to Appendix B.1 for a detailed data description.
4.3 Comparing Example Selection Methods
4.3.1 Output Inconsistency-based Methods
To evaluate the validity of inconsistency-defined uncertainty and compare the informativeness of uncertainty versus certainty instances, we implement two types of inconsistency measurement methods: Verification-based and Sampling-based methods, with $q=3$ to align with the Unc-TTP tripartite testing paradigm.
Verification-based Methods
These two-stage methods measure consistency between the LLM’s answer and its self-verification, checking if the model consistently supports its initial response. We adopt SelfCheckGPT Manakul et al. (2023) and P(True) Kadavath et al. (2022), both of which are usually used in fact-checking tasks. SelfCheckGPT scores each instance based on the consistency across $q$ rounds of verification. The higher scores indicate greater uncertainty (denoted as UncS), while lower scores reflect higher certainty (denoted as CerS). P(True) calculates the probability that the LLM outputs "True" during the verification stage, where lower probabilities indicate uncertainty, and higher probabilities suggest certainty.
Sampling-based Methods
These assess output inconsistency through multiple sampling iterations, with Unc-TTP falling into this category. The Vanilla Sampling (Vanilla) method, also known as temperature sampling, inspired by self-consistency Wang et al. (2023b), uses $q$ decoding paths followed by majority voting to identify instances of highest certainty. We extend this approach by classifying output from the $q$ samplings into the categories of CerR, CerW, and Unc.
4.3.2 Traditional Methods
To evaluate the effect of our proposed Unc-TTP-guided uncertainty as an active in-context example selection strategy, we compare our method with the Random baseline as well as other common AL approaches, following Margatina et al. (2023).
Similarity Liu et al. (2021) selects the top-$k$ semantically similar examples as demonstrations for each test sample. BM25 Robertson and Zaragoza (2009) ranks training examples based on their relevance to each test instance, and the top-ranked examples are selected as ICL demonstrations. Diversity Margatina et al. (2023) encodes all train data with Sentence-BERT Reimers and Gurevych (2019) and performs $k$-means clustering. We select one instance from each cluster. Perplexity and Entropy are two classical measures to evaluate LLMs’ uncertainty properties Margatina et al. (2023).
4.4 Implementation Details
We conduct the example selection exclusively on the training set, ensuring that we remain blind to both the validation and test sets at this stage. After that, we evaluate these candidates on the validation set to determine which uncertainty category performs best and adopt the best-performing category example for ICL testing. All ICL experiments are repeated three times with different random seeds, and we report the average accuracy along with the standard deviation. As for the prompt for evaluating the baseline method, we use the same as the one with no-label setting.
5 Experimental Results and Analysis
First, we conduct experiments with several methods that measure output inconsistency to demonstrate that uncertain instances are more informative than the certain ones when used as in-context examples. Subsequently, we compare our Unc-TTP-guided uncertainty sampling strategy with other common baselines, to further illustrate the validity of our method.
5.1 Preliminary Study: Uncertain Versus Certain Instances
Methods Cat. Llama-2 Mistral GPT-3.5
Verification-Based
P(True) CerP 66.6 (1.1) 77.5 (1.2) 81.4 (1.2)
UncP 67.0 (1.0) 76.4 (0.8) 84.4 (1.3)
SelfCheck CerS 66.7 (1.7) 77.5 (2.4) 83.8 (1.3)
GPT UncS 67.1 (1.6) 76.4 (2.2) 81.9 (1.6)
Sampling-Based
Vanilla CerW 66.7 (1.7) 78.1 (2.9) 84.4 (1.9)
CerR 66.9 (2.0) 78.2 (3.4) 83.7 (2.0)
Sampling Unc 67.6 (1.1) 79.5 (1.2) 85.4 (2.9)
Unc-TTP CerW 69.0 (4.4) 79.5 (3.4) 82.3 (2.3)
CerR 68.1 (2.7) 75.2 (2.2) 84.6 (2.4)
Unc 70.3 (2.2) 81.5 (1.5) 85.2 (1.5)
Table 2: The average accuracy on three datasets for each model over four methods that measure uncertainty via output inconsistency.
Methods Llama-2 Mistral GPT-3.5
SH HS FP Avg. SH HS FP Avg. SH HS FP Avg.
Random Baseline
Random 69.8 (1.6) 53.5 (1.5) 78.3 (1.8) 67.2 (1.6) 71.2 (1.6) 81.6 (4.4) 84.6 (2.3) 79.1 (2.8) 83.1 (4.3) 88.8 (1.8) 83.2 (0.8) 85.0 (2.3)
Traditional Active Learning Methods
Similarity* 69.9 (0.7) 52.7 (1.5) 77.3 (0.0) 66.6 (0.7) 76.0 (0.5) 78.8 (0.6) 86.1 (0.3) 80.3 (0.5) 79.2 (0.6) 87.3 (0.6) 83.4 (0.2) 83.3 (0.5)
BM25* 68.0 (0.7) 55.4 (1.1) 74.7 (1.2) 66.0 (1.0) 72.8 (0.6) 80.0 (1.1) 85.5 (0.9) 79.4 (0.9) 74.8 (1.6) 87.2 (1.0) 83.2 (1.0) 81.7 (1.2)
Diversity 66.5 (1.0) 55.9 (0.7) 81.5 (0.6) 68.0 (0.8) 53.7 (0.9) 81.1 (1.9) 84.1 (0.6) 73.0 (1.1) 62.7 (1.6) 86.7 (0.9) 81.7 (1.1) 77.0 (1.2)
Entropy✝ 68.5 (0.8) 52.7 (0.1) 80.2 (0.9) 67.1 (0.6) 68.0 (0.4) 81.5 (0.7) 83.0 (0.7) 77.5 (0.6) - - - -
Perplexity✝ 66.3 (1.3) 54.9 (1.6) 80.5 (0.7) 67.2 (1.2) 69.0 (0.7) 79.5 (0.3) 85.2 (1.3) 77.9 (0.8) - - - -
Inconsistency-Defined Uncertainty Based Active Learning Methods
P(true) 68.2 (0.8) 51.5 (1.2) 81.2 (1.1) 67.0 (1.0) 73.2 (1.0) 80.7 (0.6) 75.2 (0.8) 76.4 (0.8) 77.8 (1.0) 92.2 (0.6) 83.3 (2.2) 84.4 (1.3)
SelfCheckGPT 66.2 (1.7) 57.3 (1.9) 77.8 (1.1) 67.1 (1.6) 65.5 (2.9) 80.8 (2.1) 82.9 (1.5) 76.4 (2.2) 77.3 (3.5) 87.5 (0.7) 81.0 (0.7) 81.9 (1.6)
Vanilla Sampling 69.1 (1.0) 53.8 (1.0) 79.8 (1.2) 67.6 (1.1) 72.0 (0.9) 79.5 (1.4) 87.0 (1.2) 79.5 (1.2) 80.1 (5.5) 92.5 (1.8) 83.5 (1.3) 85.4 (2.9)
Unc-TTP 71.2 (2.1) 58.8 (3.2) 80.8 (1.3) 70.3 (2.2) 74.7 (0.8) 81.8 (2.9) 88.1 (0.8) 81.5 (1.5) 77.8 (2.4) 93.0 (0.5) 84.8 (1.7) 85.2 (1.5)
Table 3: 1-shot results of LLMs on three subjective tasks in the test set. The table presents the mean accuracy across three estimations, with standard deviations in parentheses. The highest accuracy values under each method are underlined. The best results for each dataset are highlighted in bold. Please note that * indicates methods that select different examples for each test instance, while ✝ signifies methods that require white-box access, which is not compatible with the closed-source LLM GPT-3.5.
To justify the uncertainty defined by inconsistency, we draw on the concept of uncertainty sampling in AL and apply inconsistency-defined uncertainty on in-context example selection to evaluate its effect. The results are shown in Table 2 and visualized in Figure 4.
Verification-based methods try to measure the consistency between the LLM’s first stage answer and the second stage self-verification to reveal LLM uncertainty. In this case, risking more possibility of sycophancy and cannot reflect the intrinsic uncertainty of LLM. Therefore showing no clear pattern of the relative improvement in accuracy of uncertainty examples compared to certainty examples in ICL for each model.
However, sampling-based methods, vanilla sampling and our Unc-TTP, select uncertainty instances which bring consist performance gain compare to the certain ones in most of the cases, showcasing that the selected uncertain instances are indeed more informative than the certain one under the context of ICL. Moreover, compared to vanilla, the relative performance gain is more significant for the uncertain example classified by our proposed Unc-TTP approach, which introduces label injection to strengthen the sampling rigor. As for the exception, SH on GPT-3.5, we will discuss it on the following section.
5.2 Main Results: Comparing Unc-TTP with Previous AL Methods
We have shown that the uncertainty examples selected through our Unc-TTP are more informative than the certainty examples on ICL. In this section we continue to explore the effectiveness of our proposed Unc-TTP compared to other AL strategies. The results are shown in Table 3.
For applying the traditional AL strategy on ICL, Similarity have demonstrated a obvious advantage over other strategy on GPT-3.5 and Mistral, even with the more common strategy like Entropy in traditional AL, which align with the previous results of Margatina et al. (2023). It suggests that the more capable models rely more on sample similarity while weaker model like Llama-2 depends more on example Diversity. Yet, we interestingly discover, compare to other traditional AL strategies, Random are the most informative on the strongest model GPT-3.5. It indicates that strong LLM may be insensitive to examples.
For inconsistency-defined uncertainty as the active in-context example selection strategy, ours Unc-TTP demonstrate a competitive advantage. Moreover, Unc-TTP not only performs better than the AL baseline Random by the average of 3.1%, 2.4% and 0.2% on Llama-2, Mistral and GPT-3.5, but also far exceeds the best strategy Similarity stated in previous work Margatina et al. (2023) with the improvement of 3.7%, 1.2% and 1.9%, demonstrating the effectiveness of our Unc-TTP on selecting the most informative in-context examples by the inconsistency-defined uncertainty.
5.3 Assessing the Scalability of Unc-TTP
Figure 5: Accuracy scores of Mistral $K$-way $N$-shot experiments on the test set. Our Unc-TTP robustly surpasses the random baseline and the sampling-based method when the shot number is small.
To explore the informativeness of the selected inconsistency-defined uncertain in-context examples as the number of shot increases, we conduct $N$-shot experiments on Mistral, taking the AL baseline method Random as comparison.
The results, depicted in Figure 5, reveal that uncertainty examples identified by Unc-TTP are generally more informative than those selected by the vanilla sampling-based method or chosen randomly. Unc-TTP consistently outperforms other baselines across 1 to 4 shots, highlighting the informativeness of the examples selected under Unc-TTP’s guidance. Although the advantages of inconsistency-defined uncertainty sampling diminish for the SH and FP datasets when the number of shots exceeds 8, we argue that this result is reasonable because as the number of shots increases, it is impossible to avoid introducing other variables, such as diversity and length of the instances, and the current exploration of the best combination of in-context examples is still ongoing, which will also be the direction of our future work.
5.4 Assessing the Transferability of Unc-TTP
We have shown that the wavering behavior of LLMs under limited sampling reflects their intrinsic uncertainty, and we leverage this inconsistency-defined uncertainty for active in-context example selection. In Section 3.1, we observed that more capable LLMs better resist skepticism or incorrect information, maintaining their stance while achieving higher accuracy. Based on this, we explore whether Unc-TTP selects more informative uncertain instances—where stronger LLMs waver—that provide stronger guidance for weaker LLMs compared to their self-guided instances.
[…]
Method SH HS FP Average
Llama-2
Random 69.8 (1.6) 53.5 (1.5) 78.3 (1.8) 67.2 (1.6)
UncGPT3.5 67.8 (1.5) 58.8 (1.6) 81.2 (2.2) 69.3 (1.8)
UncGPT4 71.0 (2.1) 56.8 (2.6) 80.7 (1.5) 69.5 (2.1)
UncLlama2 71.2 (2.1) 58.8 (3.2) 80.8 (1.3) 70.3 (2.2)
Mistral
Random 71.2 (1.6) 81.6 (4.4) 84.6 (2.3) 79.1 (2.8)
UncGPT3.5 71.7 (3.6) 79.6 (1.6) 86.9 (2.0) 79.4 (3.3)
UncGPT4 75.0 (2.0) 77.1 (4.9) 85.8 (2.4) 79.3 (3.1)
UncMistral 74.7 (0.8) 81.8 (2.9) 88.1 (0.8) 81.5 (1.5)
Table 4: Transferability of Unc-TTP-guided uncertainty sampling in 1-shot ICL: Comparing the informativeness of uncertainty examples selected by strong models versus those selected through weak model self-guidance.
6 Conclusion
In this paper, we introduce the Unc-TTP, a novel method for assessing intrinsic uncertainty in both open- and closed-source LLMs. From the perspective of applying active learning principles to in-context example selection, we experimentally demonstrate that uncertainty examples selected through Unc-TTP are more informative than certainty examples. This provides further evidence that output inconsistency can serve as an indicator of LLMs’ intrinsic uncertainty.
Additionally, we propose an uncertainty-based active in-context example selection strategy grounded in Unc-TTP, which outperforms previous AL strategies in terms of performance gains in ICL. These findings establish Unc-TTP as an effective tool for classifying LLM uncertainty and enhancing ICL performance, paving the way for understanding the uncertainty behavior of LLMs.
Limitations
Firstly, our proposed Unc-TTP inherits the prompt-sensitive nature of the LLMs, which may pose a risk when applying Unc-TTP-guided uncertainty sampling. However, for the robustness of our experiments, we repeated our experiments three times and reported the mean and standard deviation.
Secondly, we have only evaluated Unc-TTP on subjective tasks. Given highly subjective tasks often require a deep understanding of contextual subtleties, their labels are sometimes "biased" themselves depending on the annotator decomposition (Li et al., 2022, 2023). This characteristic is more compatible with Unc-TTP, as it can effectively expose areas where LLMs exhibit uncertainty compared to objective reasoning tasks. For the implementation of Unc-TTP on objective tasks will be a primary focus in our future work.
Thirdly, we have only tested Unc-TTP on classification tasks, as designing intuitive perturbations, like label injection, for generative or more complex tasks is challenging. This is because such perturbations inevitably introduce additional variables. One possible approach is to transform the LLM’s output into a binary fact-checking task (e.g., SelfCheckGPT). However, our paper shows that this verification-based method is ineffective under the scenario of in-context example selection. Furthermore, Unc-TTP is specifically designed for active context example selection, whereas complex generation tasks and ICL may not be compatible. Developing appropriate perturbations to extend Unc-TTP to more general generative tasks will be the focus of our future work.
Finally, although Unc-TTP adopts a one-example-for-all approach, which makes inference more efficient compared to retrieval-based methods like similarity, it requires more resources for sample selection. However, it is important to note that Unc-TTP not only serves as a novel active example selection method but also as an uncertainty classification method that can reflect the LLM’s intrinsic uncertainty. In the future, we aim to explore additional application scenarios for Unc-TTP, a perturbation-based, inconsistency-defined uncertainty classification method.
Ethics Statement
Task Time (min)
Unc-TTP on Train Set 30
1-shot ICL on Validation Set 100
1-shot ICL on Test Set 10
Table 5: The approximate computation time of the experiments. Unc-TTP on Train Set represents the classification of uncertainty categories using Unc-TTP. 1-shot ICL on Validation Set represents the evaluation of 1-shot examples from all Unc-TTP categories. 1-shot ICL on Test Set represents the application of selected uncertainty examples on the test set.
Computational Budget
We run all our experiments on Intel® Xeon® Silver 4310 CPU and A40 with a batch size of 1 and the max token number of 20. The Table 5 shows the approximate average computation time with the 7B model Llama-2 and Mistral as the LLM, using 1 random seed.
Use of AI Assistants
We have employed ChatGPT as a writing assistant, primarily for polishing the text after the initial composition.
Appendix A Detailed Performance of LLMs under Preliminary Test
The results of preliminary test under three label injection settings on three benchmarks are shown in Figure 2. LLM shows more compliance with correct labels. For the setting that provides labels, all models except Llama-2 on the SH dataset show gullibility to the given labels. Particularly, in the right-label setting, some models nearly reach 100% agreement with the given label. However, for the wrong-label setting, the model exhibits varying levels of resistance to enticement. Besides, the more capable the model is, the better it maintains accuracy in the wrong-label setting, although it does not converge to the accuracy achieved in the no-label setting. This indicates that unless the LLM is very certain of its answer, it reacts quite prudently when the labels provided deviate considerably from its own answers, as adopting the user’s point of view seems to be a safer option.
LLMs tend to make compromises when conservative options are offered, choosing the neutral option to accommodate the user’s preference. In the FP dataset, all models exhibit relatively steady performance under the wrong-label setting. A possible explanation for this observation is that FP is a classification dataset with three labels, where the distinctions between neutral and positive or negative are subtle. LLMs inherently tend to favor the more conservative neutral answers. When their initial beliefs are challenged in the wrong-label setting, they are inclined to prefer the neutral option, potentially aligning with user preferences. This behavior corresponds to the data distribution in Table 6 of the Appendix. We observe that most LLMs exhibit a tendency to produce conservative answers in the no-label setting. This tendency becomes pronounced in the wrong-label setting, where the number of neutral answers doubles, further substantiating our conclusion.
Positive Neutral Negative Failed
no-label
GPT-4 78 133 89 0
GPT-3.5 89 138 73 0
Mistral 41 179 80 0
Llama-2 120 6 174 0
right-label
GPT-4 99 101 100 0
GPT-3.5 97 110 93 0
Mistral 84 142 92 0
Llama-2 188 9 103 0
wrong-label
GPT-4 17 249 34 0
GPT-3.5 42 205 53 0
Mistral 0 280 2 18
Llama-2 82 43 175 0
Table 6: Answer distribution of no-label and wrong-label setting on Financial Phrasebank (FP), where Failed denotes the number of cases LLM fails or rejects to answer. It can be observed that the wrong-label doubles the number of neutral responses compared to the no-label setting.
Appendix B Experiment Details
B.1 Dataset Implementation Details
Since the SH dataset does not provide official data splits, we randomly selected 500 samples as the training set, 1500 as the validation set, and 200 as the test set balancing budget and time constraints.
For the FP dataset, each sentence is annotated between 5 to 8 times. To ensure high data quality, we included only those sentences with 100% annotation agreement, resulting in 2,264 examples available for training and evaluation. Among these, only 303 examples were labeled as negative. Consequently, we randomly selected 100 instances for each label to form the training set and 200 instances for each label for the validation set. We randomly sample 200 instances from the split that reach 75% annotation agreement for the test set. In the wrong-label setting of Unc-TTP, for datasets with more than two labels, such as FP, we randomly select an incorrect label to serve as the provided label.
Other data splits are detailed in Table 7 for reference. It should be noted that all divisions have an equal number of examples for each label. The labels we use in the prompt are as follows:
  • SH: sarcastic / non-sarcastic
  • HS: humorous / not humorous
  • FP: positive / neutral / negative
  Dataset Train Valid Test Label
SH 500 1500 200 2
HS 500 1500 200 2
FP 300 600 300 3
Table 7: The amount of data and the corresponding number of labels for each dataset we used in our experiments. For all splits, the number of labels is equally distributed.
B.2 Example Selection Details
Uncertainty Category Distribution
Figure 6: Number of examples corresponding to the respective labels for each model on the three datasets.
We visualize the respective quantities of instances for each category in Figure 6. Interestingly, the number of instances in the categories that do not match human intuition, such as 001, 100, and 101, is notably low, as expected. For instance, if a model misclassifies an instance both without guidance and with correct guidance, it should have a very low likelihood of answering correctly when provided with incorrect labels.
Moreover, the categories we anticipate being the most informative—those that could stimulate deep thinking in the model when presented with a label, such as 011—are also scarce. This scarcity is likely due to the high level of capability required to generate such categories. The next most informative categories, in our view, are 010 and 110. These three categories represent the uncertainty types of primary interest to us.
Appendix C Uncertainty Category Selection
We select the final inconsistency-defined uncertainty category of both Unc-TTP and vanilla sampling using validation set due to the category where we regard as informative, such as 011 is scarce, and sometime have not enough data in the particular category to perform $K$-way $N$-shot experiment.
The Unc-TTP-guided uncertainty sampling result on validation set are shown in the Table 9. For smaller models, the informative uncertainty categories are predominantly centered around 011 and 010. However, for larger models, there is no discernible pattern in the informative uncertainty categories. We speculate that this lack of pattern may be due to two reasons: first, the larger model may rely less on in-context learning (ICL) compared to smaller models; and second, the increased parameter size of the larger model introduces greater stochasticity, even when using greedy decoding.
Type SH HS FP
Llama-2
000 63.9 (2.3) 52.8 (1.8) 85.5 (0.8)
111 63.9 (2.9) 52.9 (1.4) 87.5 (1.2)
001/010/100 62.5 (0.6) 53.8 (1.1) 82.1 (3.7)
011/101/110 64.6 (0.9) 53.5 (3.6) 84.6 (2.8)
Mistral
000 65.7 (5.1) 80.3 (3.1) 89.2 (2.3)
111 71.2 (3.8) 76.2 (0.7) 90.0 (0.7)
001/010/100 72.0 (0.9) 76.5 (2.1) 90.4 (0.9)
011/101/110 69.2 (4.6) 81.2 (1.1) 89.3 (3.6)
GPT-3.5
000 71.0 (2.9) 91.9 (0.5) 91.0 (1.6)
111 76.2 (3.3) 91.1 (1.2) 90.5 (1.7)
001/010/100 73.4 (1.8) 91.8 (1.9) 89.8 (2.1)
011/101/110 77.6 (3.1) 90.9 (0.1) 88.7 (2.8)
Table 8: Vanilla Sampling method on validation set.
The utility of Vanilla-Sampling-selected uncertainty examples for each category in the validation set is shown in Table 8. In vanilla sampling, the number of certainty and uncertainty categories is identical, with each category consisting of two types. For the certainty category, the types are "certainly wrong" (000) and "certainly right" (111). For the uncertainty category, the types are "one right, two wrong" and "one wrong, two right." Therefore, the selection of uncertainty examples is completely balanced. The performance gain observed when uncertainty examples are used in ICL, compared to when certainty examples are used, is not due to any bias in validation set selection.
Type SH HS FP
Llama-2
000 65.5 (1.6) 60.3 (5.8) 86.2 (0.4)
111 65.7 (3.1) 55.6 (2.9) 83.8 (2.3)
001 64.4 (1.5) - 84.9 (1.6)
010 68.6 (0.6) 58.2 (2.8) 84.9 (2.6)
011 68.0 (1.2) 62.4 (1.8) 86.5 (1.9)
100 - - 84.6 (0.9)
101 66.5 (2.4) - 86.6 (1.4)
110 - 57.0 (4.1) 84.1 (0.9)
Random 66.8 (1.7) 53.4 (3.3) 84.4 (3.2)
Mistral
000 68.2 (3.6) 81.9 (2.0) 90.6 (4.0)
111 65.4 (2.9) 75.3 (0.6) 88.9 (2.1)
001 - - -
010 69.6 (2.0) 81.4 (0.9) 91.1 (1.0)
011 - 82.0 (2.6) 86.3 (0.9)
100 - - -
101 - - -
110 68.6 (1.9) 80.9 (2.9) 89.9 (1.9)
Random 67.3 (1.2) 81.8 (3.2) 89.2 (4.3)
GPT-3.5
000 77.0 (2.1) 88.4 (1.4) 83.9 (2.1)
111 81.6 (2.9) 89.7 (0.8) 87.2 (1.5)
001 - - -
010 77.0 (3.8) 88.4 (1.5) 85.3 (1.7)
011 - 89.8 (1.0) 85.6 (2.5)
100 - 92.3 (0.3) 86.9 (1.1)
101 - 90.0 (2.3) -
110 76.6 (4.3) 90.0 (1.4) 88.3 (2.8)
Random 80.4 (2.2) 89.8 (1.4) 86.2 (2.6)
Table 9: 1-shot results of LLMs in the validation set. The table presents the mean accuracy across three estimations, with standard deviations in parentheses. The two highest accuracy values under the Unc-TTP setting are underlined. The highest results are highlighted in bold.
Appendix D Detailed Results on Accessing the Informative of Inconsistency-Based Methods
We present the complete results on three datasets for each of the four methods of measuring uncertainty through output inconsistency in Figure 10.
Methods Cat. SH HS FP Average
Llama-2
P(True) Max 70.0 (0.4) 53.6 (1.1) 76.1 (1.8) 66.6 (1.1)
Min 68.2 (0.8) 51.5 (1.2) 81.2 (1.1) 67.0 (1.0)
SelfCheck CerS 66.0 (2.5) 56.7 (1.2) 77.3 (1.3) 66.7 (1.7)
GPT UncS 66.2 (1.7) 57.3 (1.9) 77.8 (1.1) 67.1 (1.6)
Sampling CerW 67.9 (2.7) 53.3 (2.1) 78.8 (0.2) 66.7 (1.7)
CerR 66.9 (2.4) 53.8 (1.8) 80.0 (1.9) 66.9 (2.0)
Unc 69.1 (1.0) 53.8 (1.0) 79.8 (1.2) 67.6 (1.1)
Unc-TTP (ours) CerW 68.5 (2.0) 59.5 (9.5) 79.1 (1.7) 69.0 (4.4)
CerR 70.4 (1.9) 56.9 (2.9) 77.0 (3.4) 68.1 (2.7)
Unc 71.2 (2.1) 58.8 (3.2) 80.8 (1.3) 70.3 (2.2)
Mistral
P(True) Max 70.3 (0.6) 80.7 (2.3) 81.5 (0.8) 77.5 (1.2)
Min 73.2 (1.0) 80.7 (0.6) 75.2 (0.8) 76.4 (0.8)
SelfCheck CerS 70.2 (3.0) 82.0 (1.4) 80.2 (2.9) 77.5 (2.4)
GPT UncS 65.5 (2.9) 80.8 (2.1) 82.9 (1.5) 76.4 (2.2)
Sampling CerW 65.7 (5.1) 81.3 (1.8) 87.3 (1.9) 78.1 (2.9)
CerR 71.2 (3.8) 77.3 (5.0) 86.0 (1.4) 78.2 (3.4)
Unc 72.0 (0.9) 79.5 (1.4) 87.0 (1.2) 79.5 (1.2)
Unc-TTP (ours) CerW 71.2 (5.4) 81.0 (2.3) 86.4 (2.4) 79.5 (3.4)
CerR 66.6 (3.8) 74.6 (2.0) 84.3 (0.9) 75.2 (2.2)
Unc 74.7 (0.8) 81.8 (2.9) 88.1 (0.8) 81.5 (1.5)
GPT-3.5
P(True) Max 74.8 (1.7) 89.3 (0.9) 80.2 (1.1) 81.4 (1.2)
Min 77.8 (1.0) 92.2 (0.6) 83.3 (2.2) 84.4 (1.3)
SelfCheck CerS 78.7 (2.9) 89.8 (0.5) 82.9 (0.6) 83.8 (1.3)
GPT UncS 77.3 (3.5) 87.5 (0.7) 81.0 (0.7) 81.9 (1.6)
Sampling CerW 78.5 (3.9) 91.7 (0.8) 83.0 (1.1) 84.4 (1.9)
CerR 77.3 (2.9) 91.0 (1.3) 82.7 (1.8) 83.7 (2.0)
Unc 80.1 (5.5) 92.5 (1.8) 83.5 (1.3) 85.4 (2.9)
Unc-TTP (ours) CerW 78.9 (3.6) 88.2 (1.8) 79.8 (1.4) 82.3 (2.3)
CerR 82.8 (2.9) 88.8 (2.4) 82.1 (2.0) 84.6 (2.4)
Unc 77.8 (2.4) 93.0 (0.5) 84.8 (1.7) 85.2 (1.5)
Table 10: Complete results on three datasets for each of the four methods of measuring uncertainty through output inconsistency for each model.